OpenStudy (anonymous):

The sum of two numbers is 52. The greater number is 4 more than the smaller number. Which equation can be used to solve for the smaller number?

OpenStudy (anonymous):

@jabez177 @briannabug @AloneS @cutegirl

jhonyy9 (jhonyy9):

x+y=52 y=4+x

OpenStudy (anonymous):

x − (x + 4) = 52 x + (x + 4) = 52 x(x + 4) = 52 x(x − 4) = 52